Using divisibility tests, determine which of the following numbers are divisible by 11: 
5445 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Sum of the digits at odd places = 5 + 4 =9
Sum of the digits at even places = 4 + 5 =9
Difference = 9–9 = 0
Since, the difference between sum of digits at odd places and sum of digits at even places is 0.
Hence,5445 is divisible by 11

Thanks